OpenStudy (anonymous):

me and my twin sister are arguing about this problem can someone tell us the steps to this please.

OpenStudy (anonymous):

simplest radical

Nnesha (nnesha):

\[\huge\rm \frac{ 1 }{ -\frac{ 3 }{ 8 } }\] is it ur question ?? :o ;

OpenStudy (anonymous):

yesss :)

Nnesha (nnesha):

waoo so multiply numerator by the reciprocal of the denominator fraction

OpenStudy (anonymous):

okhay have the steps in order so she may understande

OpenStudy (anonymous):

understand

Nnesha (nnesha):

u sure u posted the right question ?

Nnesha (nnesha):

example \[ \huge\rm \frac{ a }{ \frac{ b }{ c} } = a \times \frac{ c }{ b }\] multiply top withe the reciprocal of the bottom fraction that's ti

OpenStudy (anonymous):



Nnesha (nnesha):

what is that ?